OpenStudy (anonymous):

Simplify the expression, i^19 and i^97.

OpenStudy (chaise):

i^1 = sqrt(-1) i^2 = -1 i^3 = -i i^4 = 1 This repeats again, so i^5 = sqrt(-1). So every 4 it repeats. So i^16 = 1. i^17 = sqrt(-1) i^18 = -1 i^19 = -i Does this make sense? Can you do 96 on your own?

OpenStudy (chaise):

97*

OpenStudy (anonymous):

I know how to do it. I was just lazy and didn't want to think about it. hah. Sorry.

OpenStudy (chaise):

I will do it on my calculator. :) i^97 = i

OpenStudy (anonymous):

Thank you! :)